# Backtracking New Q-Newton’s method for finding roots of meromorphic functions in 1 complex variable: Global convergence, and local stable/unstable curves

### v1

@@ 4.4. Some (both rigorous and heuristic) explanations for the observed experiments on BNQN
With these choices and assume that $f$ is a polynomial, Theorem 1.3 can be used to rigorously explain the observations, provided one can show that for $z^{*}\in\mathcal{C}(f)$ and the disc $\mathbb{D}(z^{*},r)$ for which Theorem 1.4 applies, then $\mathcal{E}\cap\mathbb{D}(z^{*},r)$ is contained in a finite union of real analytic curves. From Theorem 1.4, we know that the $d$ stable curves are real analytic and belong to $\mathcal{E}$. Hence, the difficulty lies at the fact that there could be ”bad” initial points $z_{0}\in\mathbb{D}(z^{*},r)$ for which some $z_{k}$ leaves the disc, but then later some $z_{k^{\prime}}$ re-enters the disc and lands on the stable curves. If one can show, for a given function $f$, that such ”bad” initial points do not exist, then one rigorously establishes that $\mathcal{E}$ is - away from poles of $f$ contained in a finite union of real analytic curves.
Observation 2: BNQN seems similar to Newton’s flow.
This can be explained as follows. By Theorem 1.4, near any critical point $z^{*}$ but not a root of $f$ (which is allowed to be any meromorphic function), and of order $d$, there are precisely $d$ stable curves for the dynamics of Analytic BNQN. For Newton’s flow, there is a similar fact, see e.g. the discussion on critical lines in [11, Section 5] for the case of rational functions.
@@
Figure 6. Basins of attraction for finding the roots of $f(z)=1-z^{4}$. Top picture: Voronoi’s diagram of the 4 roots, center picture: Newton’s method, bottom picture: BNQN. Points of the same colour belongs to basin of attraction of the same root. Points in black do not belong to the basin of attraction of any of the concerned 4 roots. Here the parameter $\theta$ in Algorithm 1 is $1$.
@@
Figure 7. Basins of attraction for finding the roots of $f(z)=(z-1)(z-2)(z+1)(z-7)$. Top picture: Voronoi’s diagram of the 4 roots, center picture: Newton’s method, bottom picture: BNQN. Points of the same colour belongs to basin of attraction of the same root. Points in black do not belong to the basin of attraction of any of the concerned 4 roots. Here the parameter $\theta$ in Algorithm 1 is $1$.
@@
Figure 8. Basins of attraction for finding the roots of $f(z)=(z+2)(z-5)(z-1-4i)z$. Top picture: Voronoi’s diagram of the 4 roots, center picture: Newton’s method, bottom picture: BNQN. Points of the same colour belongs to basin of attraction of the same root. Points in black do not belong to the basin of attraction of any of the concerned 4 roots. Here the parameter $\theta$ in Algorithm 1 is $1$.
@@
Figure 9. Basins of attraction for finding the roots of $f(z)=(z+2)(z-5)(z-1-4i)(z-3-i)$. Top picture: Voronoi’s diagram of the 4 roots, center picture: Newton’s method, bottom picture: BNQN. Points of the same colour belongs to basin of attraction of the same root. Points in black do not belong to the basin of attraction of any of the concerned 4 roots. Here the parameter $\theta$ in Algorithm 1 is $1$.
@@
Figure 10. Basins of attraction for finding the roots of $f(z)=P(z)/P^{\prime}(z)$, where $P(z)=z(z-2i)(z-5-2i)(z-3+3i)(z-2-i)$. Top picture: Voronoi’s diagram of the 5 roots, center picture (from [9]): Newton’s method, bottom picture: BNQN. Points of the same colour belongs to basin of attraction of the same root. Points in black do not belong to the basin of attraction of any of the concerned 5 roots. Here the parameter $\theta$ in Algorithm 1 is $1$: note that the picture for BNQN here looks more smooth than that in [9], where $\theta=0$ was used.

### v2

@@ 4.4. Some (both rigorous and heuristic) explanations for the observed experiments on BNQN
With these choices and assume that $f$ is a polynomial (or more generally has compact sublevels), Theorem 1.3 can be used to rigorously explain the observations, provided one can show that for $z^{*}\in\mathcal{C}(f)$ and the disc $\mathbb{D}(z^{*},r)$ for which Theorem 1.4 applies, then $\mathcal{E}\cap\mathbb{D}(z^{*},r)$ is contained in a finite union of real analytic curves. From Theorem 1.4, we know that the $d$ stable curves are real analytic and belong to $\mathcal{E}$. Hence, the difficulty lies at the fact that there could be ”bad” initial points $z_{0}\in\mathbb{D}(z^{*},r)$ for which some $z_{k}$ leaves the disc, but then later some $z_{k^{\prime}}$ re-enters the disc and lands on the stable curves. If one can show, for a given function $f$, that such ”bad” initial points do not exist, then one rigorously establishes that $\mathcal{E}$ is - away from poles of $f$ contained in a finite union of real analytic curves.
Observation 2: BNQN behaviour seems similar to Newton’s flow and Poincaré-Bendixon theorem.
We recall that Newton’s flow method for finding roots of a meromorphic function $f$ is to find solutions $z(t)$ of the following ODE:
$\frac{dz(t)}{dt}=-\frac{f(z(t))}{f^{\prime}(z(t))},$
with initial value $z(0)=z_{0}$. One would like to show that $\lim_{t\rightarrow\infty}z(t)$ is a root of $f(z)$.
In [10], it has been found by many experiments that the basins of attraction for BNQN and Newton’s method look very similar. Here, we will provide an extensive discussion on this similarity, based on the results in this paper.
Usually, the dynamics of an ODE is better behaved than its discrete version. In our case, that Newton’s flow (the ODE) behaves better than its discrete version (Relaxed Newton’s method) is already known since the 1970s. Here, we recall some main facts about the behaviour of Newton’s flow for finding roots of a meromorphic function $f$, following [9] (the reader can find more detail in this paper and the survey [5], as well as references therein), and compare with BNQN.
An important property of Newton’s flow is that: $f(z(t))=e^{-t}f(z_{0})$ (in the open interval - could be bounded - where the solution exists). Based on this, even originally the right hand side of the ODE for Newton’s flow (i.e. $-f(z)/f^{\prime}(z)$) is not real analytic (even not continuous), one can multiply it with a non-negative function (more precisely, $-(1+|f(z)|^{4})^{-1}|f(z)|^{2}$) to obtain a new ODE:
(4.1) $\frac{dz(t)}{dt}=g(z(t))$
where $g(z)$ is a real analytic function, and solutions of this ODE and Newton’s flow have the same trajectories. Therefore, studying the limit behaviour of Newton’s flow is the same as studying the limit behaviour of the new ODE.
Based on this, one can prove many facts, for example: A limit point of Newton’s flow is either a root of $f$, a critical point of $f$, or $\infty$. This fact is known also for BNQN, see [28].
Moreover, since $g(z(t))$ is real analytic, the classical Poincaré-Bendixon theorem applies to the ODE (4.1). In particular, it follows that the behaviour of Newton’s flow near a critical point of order $d$ of $f$ (but is not itself a root of $f$) is that of $d$-fold saddle points. This is exactly the same as that described in Theorem 1.4 in the current paper.
Theorem 1.2 says that the set of initial points $z_{0}$ for which BNQN does not converge to a root of $f$ or $\infty$ is locally a countable union of real analytic curves. A similar (and stronger) property is known for Newton’s method.
We see therefore that indeed there are many similarities between BNQN (an iterative method) and Newton’s flow (a continuous method). This is the more striking given that a fixed discrete version of Newton’s flow (e.g. Relaxed Newton’s method when $\gamma$ is fixed or the explicit Runger-Kutta with a fixed set of parameters) is known to not have these properties, thanks to [16]. Remark that when doing experiments in practice, one needs to use one of the discrete versions of Newton’s flow, for which it is not guaranteed (by the mentioned reason) that the pictures obtained reflect well the theoretical results for Newton’s flow. It is also worth noting that for the Relaxed Newton’s method to be a reasonable approximation of Newton’s flow, one needs to choose $\gamma$ to be small, which may make the running time to be very long (and this is actually observed in practice).
@@
Figure 8. Basins of attraction for finding the roots of $f(z)=1-z^{4}$. Top picture: Voronoi’s diagram of the 4 roots, center picture: Newton’s method, bottom picture: BNQN. Points of the same colour belong to basin of attraction of the same root. Here the parameter $\theta$ in Algorithm 1 is $1$.
@@
Figure 9. Basins of attraction for finding the roots of $f(z)=(z-1)(z-2)(z+1)(z-7)$. Top picture: Voronoi’s diagram of the 4 roots, center picture: Newton’s method, bottom picture: BNQN. Points of the same colour belong to basin of attraction of the same root. Here the parameter $\theta$ in Algorithm 1 is $1$.
@@
Figure 10. Basins of attraction for finding the roots of $f(z)=(z+2)(z-5)(z-1-4i)z$. Top picture: Voronoi’s diagram of the 4 roots, center picture: Newton’s method, bottom picture: BNQN. Points of the same colour belong to basin of attraction of the same root. Here the parameter $\theta$ in Algorithm 1 is $1$.
@@
Figure 11. Basins of attraction for finding the roots of $f(z)=(z+2)(z-5)(z-1-4i)(z-3-i)$. Top picture: Voronoi’s diagram of the 4 roots, center picture: Newton’s method, bottom picture: BNQN. Points of the same colour belong to basin of attraction of the same root. Here the parameter $\theta$ in Algorithm 1 is $1$.
@@
Figure 12. Basins of attraction for finding the roots of $f(z)=P(z)/P^{\prime}(z)$, where $P(z)=z(z-2i)(z-5-2i)(z-3+3i)(z-2-i)$. Top picture: Voronoi’s diagram of the 5 roots, center picture (from [10]): Newton’s method, bottom picture: BNQN. Points of the same colour belong to basin of attraction of the same root. Here the parameter $\theta$ in Algorithm 1 is $1$: note that the picture for BNQN here looks more smooth than that in [10], where $\theta=0$ was used.